# Patch-wise Structural Loss for Time Series Forecasting

### v1

@@ Appendix C Related Work
Appendix D Running Cost
To evaluate the computational cost introduced by PS loss, we measure the running time per iteration (ms/iter) of backbone models trained with MSE loss and PS loss across different forecasting lengths on the ETTh1 dataset. As shown in Table 7, although PS loss results in a modest increase in computation time, it consistently enhances forecasting accuracy across all models. Notably, DLinear achieves the highest average MSE improvement of 6.1%, highlighting the efficiency-accuracy trade-off introduced by PS loss.
Table 7: Running cost of PS loss. The table reports the running time per iteration (ms/iter) for backbone models using MSE loss and PS loss on the ETTh1 dataset. “Time Inc.” represents the additional running time introduced by PS loss, while “MSE Dec.” indicates the forecasting performance improvement achieved with PS loss.
Models iTransformer Dlinear TimesNet
Running Cost MSE Loss PS Loss Time Inc. MSE Dec. MSE Loss PS Loss Time Inc. MSE Dec. MSE Loss PS Loss Time Inc. MSE Dec.
(ms/iter) (ms/iter) (ms/iter) (%) (ms/iter) (ms/iter) (ms/iter) (%) (ms/iter) (ms/iter) (ms/iter) (%)
ETTh1 96 22.9 30.4 7.6 2.1% 12.5 19.3 6.9 4.4% 65.2 70.9 5.7 0.6%
192 24.7 31.3 6.6 2.9% 12.0 19.1 7.1 9.2% 66.5 74.9 8.3 4.5%
336 26.5 32.1 5.6 3.4% 12.2 19.4 7.2 2.7% 84.3 92.7 8.4 7.5%
720 27.7 31.7 4.0 5.7% 14.2 21.3 7.1 8.2% 107.5 118.4 10.9 5.4%
Avg 25.4 31.3 5.9 3.5% 12.7 19.7 7.1 6.1% 80.8 89.2 8.3 4.5%
Appendix E Impact of PS Loss on Generalization
@@
From Figure  6, we observe a consistent trend across both datasets. During training, models optimized solely with MSE loss achieve lower training error per epoch compared to those optimized with PS loss. However, on testing data, models trained with MSE loss exhibit higher testing error per epoch compared to those trained with PS loss. These observations suggest that models trained with MSE loss focus solely on the point-wise error, resulting in smaller training losses, but fail to generalize effectively to the testing data. In contrast, PS loss mitigates overfitting by encouraging the model to focus on local structural similarity in addition to point-wise accuracy, highlighting its effectiveness in improving generalization and forecasting accuracy.
Appendix F More Results on Hyperparameter Sensitivity
(a) iTransformer as backbone. (left: ETTm2, right: Weather)
(b) DLinear as backbone. (left: ETTm2, right: Weather)
Appendix G Forecasting Visualization
@@
Appendix H More Results on Ablation Study

### v2

@@ Appendix C Related Work
Appendix D Computational Cost
To evaluate the computational cost introduced by the proposed PS loss, we conducted a theoretical complexity analysis and measured the actual runtime overhead during training.
D.1 Time Complexity Analysis
We analyze the complexity of our proposed PS loss with respect to the forecasting length $T$, channel number $C$, and the last hidden layer’s dimension $d$. The overall time complexity comes from three main components:
  • Fourier-based Adaptive Patching (FAP). The time complexity for this component is dominated by the Fast Fourier Transform (FFT) operation, which has a complexity of $O(T\log T)$ for a time series of length T. Since the FFT is applied independently to each of the C channels, the total time complexity is $O(C\cdot T\log T)$.
  • Patch-wise Structural Loss (PS). The series is divided into $N=\lfloor\frac{(T-P)}{S}\rfloor+1\approx\frac{2T}{P}$ patches, where $P$ denotes the patch length and the stride $S$ is set to $\lfloor\frac{P}{2}\rfloor$. Computing the correlation, variance, and mean for each patch requires $P$ operations. With a total of $C\cdot N$ patches, the time complexity of this component is $O(C\cdot N\cdot P)=O(C\cdot T)$.
  • Gradient-based Dynamic Weighting (GDW). The gradient computation for each loss term with respect to the output layer weights, which have shape $d\times T$, requires $O(d\cdot T)$ operations. Considering all three loss components, the total time complexity of GDW is $O(3\cdot d\cdot T)=O(d\cdot T)$.
Therefore, the overall time complexity of the PS loss is $O(C\cdot TlogT+C\cdot T+d\cdot T)$.
D.2 Actual Runtime Overhead
To assess the runtime overhead introduced by the PS loss, we measure the training time per epoch (seconds/epoch) for both the standard MSE loss and the proposed PS loss, using the iTransformer model as the backbone. We select three datasets of different scales for the experiment: ETTh1 (small), Weather (medium), and ECL (large). As shown in Table 7, while PS loss introduces additional computational cost, the increase in training time remains modest across datasets of different sizes. This overhead is justified by the consistent improvements in forecasting accuracy achieved with the PS loss.
Table 7: Runtime Overhead of PS loss. The table reports the running time per epoch (seconds/epoch) for iTransformer model using MSE loss and PS loss on the three datasets: ETTh1, Weather, and ECL. “Time Inc.” represents the additional running time introduced by PS loss.
Datasets ETTh1 Weather ECL
Train Time (s/epoch) MSE Time PS Time Time Inc. MSE Time PS Time Time Inc. MSE Time PS Time Time Inc.
iTransformer 96 1.99 2.64 0.65 10.25 13.27 3.03 24.03 27.36 3.34
192 1.98 2.73 0.74 10.62 14.17 3.55 24.35 28.23 3.89
336 1.91 2.64 0.73 10.71 14.52 3.81 25.15 30.13 4.98
720 1.95 2.65 0.70 10.96 14.19 3.23 26.55 35.07 8.52
Avg. 1.96 2.66 0.71 10.63 14.04 3.40 25.02 30.20 5.18
Appendix E Additional Evaluation Metrics for PS Loss Performance
To provide a more comprehensive evaluation of the PS loss, we incorporate additional metrics, including Dynamic Time Warping (DTW) (Müller, 2007), Time Distortion Index (TDI) (Frías-Paredes et al., 2016), and Pearson Correlation Coefficient (PCC) (Cohen et al., 2009), to assess the quality of the forecasting results.
E.1 Metric Definitions
Dynamic Time Warping.
DTW computes the distance between two time series by identifying an optimal warping path, allowing non-linear alignments along the time axis to emphasize shape similarity (Müller, 2007). Given two time series $Y=\{y_{0},y_{1},\dots,y_{T-1}\}\in\mathbb{R}^{T}$ and $\hat{Y}=\{\hat{y}_{0},\hat{y}_{1},\dots,\hat{y}_{T-1}\}\in\mathbb{R}^{T}$, the DTW distance is defined as:
$\mathrm{DTW}(Y,\hat{Y})=\min_{\mathbf{A}\in\mathcal{A}(Y,\hat{Y})}\sum_{(i,j)\in\mathbf{A}}d(y_{i},\hat{y}_{j})=\sum_{(i,j)\in\mathbf{A}^{*}}d(y_{i},\hat{y}_{j}),$
where $d(\cdot,\cdot)$ denotes a distance measure, typically the squared Euclidean distance. $\mathbf{A}$ is a warping path consisting of $K$ index pairs $\{(i_{0},j_{0}),(i_{1},j_{1}),\dots,(i_{K-1},j_{K-1})\}$, where $10\leq i_{k},j_{k}\leq T-$ 1. $\mathcal{A}(Y,\hat{Y})$ denotes the set of all admissible warping paths, and $\mathbf{A}^{*}\in\mathcal{A}(Y,\hat{Y})$ is the optimal path that yields the minimum total distance between aligned time steps. A warping path $\mathbf{A}$ is considered admissible if it satisfies the following conditions:
  • Boundary condition: $(i_{0},j_{0})=(0,0)$ and $(i_{K-1},j_{K-1})=(T-1,T-1)$
  • Monotonicity condition: $i_{k+1}\geq i_{k}$ and $j_{k+1}\geq j_{k}$, for all $k\in[0,K-2]$
  • Step size condition: $(i_{k+1}-i_{k},j_{k+1}-j_{k})\in\{(1,0),(0,1),(1,1)\}$, for all $k\in[0,K-2]$
Time Distortion Index.
TDI quantifies the degree of temporal distortion between two sequences (Frías-Paredes et al., 2016). It was originally defined as the area between the optimal DTW warping path $\mathbf{A}^{*}$ and the identity path =$\mathcal{I}=\{(t,t)\}_{t=0}^{T-1}$. In our work, we adopt a simplified version of TDI proposed in (Le Guen & Thome, 2019). Given two time series $Y\in\mathbb{R}^{T}$ and $\hat{Y}\in\mathbb{R}^{T}$ and their optimal warping path $\mathbf{A}^{*}$, the TDI is calculated as:
$\mathrm{TDI}(Y,\hat{Y})=\sum_{(i,j)\in\mathbf{A}^{*}}\frac{(i-j)^{2}}{T^{2}}.$
Pearson Correlation Coefficient.
PCC measures the degree of linear correlation between two time series. Given $Y=\{y_{0},y_{1},\dots,y_{T-1}\}\in\mathbb{R}^{T}$ and $\hat{Y}=\{\hat{y}_{0},\hat{y}_{1},\dots,\hat{y}_{T-1}\}\in\mathbb{R}^{T}$, the PCC is calculated as:
$\text{PCC}(Y,\hat{Y})=\frac{\sum_{t=0}^{T-1}(y_{t}-\bar{y})(\hat{y}_{t}-\bar{\hat{y}})}{\sqrt{\sum_{t=0}^{T-1}(y_{t}-\bar{y})^{2}}\cdot\sqrt{\sum_{t=0}^{T-1}(\hat{y}_{t}-\bar{\hat{y}})^{2}}},$
where $\bar{y}$ and $\bar{\hat{y}}$ denote the mean values of $Y$ and $\hat{Y}$, respectively.
E.2 Evaluation Results
We evaluated the forecasting performance of the iTransformer model trained with MSE and PS loss using DTW, TDI, and PCC. The results are presented in Table 8. PS loss consistently improves all three shape-aware metrics, indicating better structural alignment. On the ETTh2 dataset, the iTransformer trained with MSE achieves lower DTW scores, reflecting smaller numerical differences after optimal alignment. However, the higher TDI in this case suggests that the alignment involves more extensive temporal warping, indicating greater structural distortion compared to the forecasts generated using PS loss.
Table 8: Additional evaluation metrics for assessing the performance of PS loss. This table reports the performance of the iTransformer model trained with MSE and PS loss, evaluated using three shape-aware metrics: DTW, TDI, and PCC.
Model iTransformer
Metrics MSE DTW TDI PCC
Loss Functions MSE + PS MSE + PS MSE + PS MSE + PS
ETTh1 96 0.387 0.379 3.743 3.725 2.475 2.278 0.562 0.572
192 0.441 0.428 5.728 5.685 4.549 4.479 0.530 0.539
336 0.491 0.474 8.055 8.042 9.356 8.393 0.496 0.519
720 0.509 0.480 11.896 11.844 15.171 12.686 0.469 0.488
Avg 0.457 0.440 7.355 7.324 7.888 6.959 0.514 0.530
ETTh2 96 0.301 0.289 3.256 3.213 5.739 5.207 0.355 0.408
192 0.380 0.373 5.227 5.321 14.124 13.937 0.311 0.353
336 0.424 0.416 7.494 7.644 27.298 24.994 0.286 0.325
720 0.430 0.421 11.586 11.886 51.733 46.683 0.244 0.283
Avg 0.384 0.375 6.891 7.016 24.723 22.705 0.299 0.342
ETTm1 96 0.342 0.326 3.275 3.167 6.199 5.569 0.575 0.604
192 0.383 0.374 4.887 4.796 8.221 7.561 0.556 0.573
336 0.418 0.410 6.917 6.815 11.863 10.831 0.530 0.545
720 0.487 0.472 11.195 10.960 23.522 21.562 0.490 0.505
Avg 0.408 0.396 6.568 6.435 12.451 11.381 0.538 0.557
ETTm2 96 0.186 0.175 2.575 2.358 6.338 5.544 0.364 0.425
192 0.254 0.242 4.141 3.906 11.684 10.748 0.337 0.403
336 0.316 0.304 6.167 5.878 25.611 21.580 0.318 0.377
720 0.414 0.401 10.770 10.300 64.242 52.106 0.281 0.341
Avg 0.292 0.281 5.913 5.611 26.969 22.495 0.325 0.387
Weather 96 0.176 0.167 2.174 2.107 11.990 11.166 0.344 0.381
192 0.227 0.219 3.725 3.703 22.741 21.400 0.339 0.371
336 0.282 0.274 5.768 5.772 42.628 41.029 0.319 0.344
720 0.357 0.353 9.975 10.053 88.401 87.777 0.296 0.313
Avg 0.261 0.253 5.410 5.409 41.440 40.343 0.324 0.352
Appendix F More Results on Hyperparameter Sensitivity
(a) iTransformer as backbone. (left: ETTm2, right: Weather)
(b) DLinear as backbone. (left: ETTm2, right: Weather)
Appendix G Impact of PS Loss on Generalization
@@
From Figure  8, we observe a consistent trend across both datasets. During training, models optimized solely with MSE loss achieve lower training error per epoch compared to those optimized with PS loss. However, on testing data, models trained with MSE loss exhibit higher testing error per epoch compared to those trained with PS loss. These observations suggest that models trained with MSE loss focus solely on the point-wise error, resulting in smaller training losses, but fail to generalize effectively to the testing data. In contrast, PS loss mitigates overfitting by encouraging the model to focus on local structural similarity in addition to point-wise accuracy, highlighting its effectiveness in improving generalization and forecasting accuracy.
Appendix H Forecasting Visualization
@@
Appendix I More Results on Ablation Study
@@
Appendix J More Experiments on Zero-shot Forecasting
Different Forecasting Lengths.
We conducted zero-shot forecasting experiments on the iTransformer model with forecasting lengths of 96, 336, and 720. As results shown in Table 11, PS loss consistently enhances forecasting accuracy across all horizons, yielding improvements in 33 out of 36 settings. These findings underscore the effectiveness of PS loss in improving zero-shot forecasting performance under varying prediction lengths.
Table 11: PS loss zero-shot performance on different forecasting lengths using iTransformer as the backbone model. The forecasting lengths are {96, 336, 720}.
Forecasting Length 96 336 720
Loss Function MSE PS MSE PS MSE PS
Source Target MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 ETTh2 0.296 0.344 0.296 0.342 0.424 0.431 0.422 0.425 0.429 0.444 0.432 0.445
ETTm1 0.963 0.611 0.821 0.570 0.845 0.597 0.949 0.616 0.866 0.610 0.828 0.601
ETTm2 0.238 0.321 0.222 0.307 0.342 0.375 0.353 0.382 0.442 0.428 0.436 0.424
ETTh2 ETTh1 0.575 0.515 0.465 0.455 0.656 0.559 0.607 0.534 0.708 0.603 0.626 0.563
ETTm1 1.066 0.641 0.927 0.597 0.918 0.615 0.887 0.603 0.906 0.629 0.869 0.614
ETTm2 0.237 0.321 0.228 0.313 0.349 0.381 0.347 0.380 0.445 0.430 0.441 0.428
ETTm1 ETTh1 0.708 0.556 0.617 0.521 0.731 0.576 0.626 0.532 0.743 0.599 0.633 0.559
ETTh2 0.351 0.390 0.342 0.380 0.490 0.469 0.473 0.457 0.486 0.476 0.467 0.463
ETTm2 0.202 0.279 0.198 0.272 0.322 0.352 0.319 0.347 0.422 0.407 0.419 0.403
ETTm2 ETTh1 0.833 0.603 0.568 0.497 1.119 0.716 0.642 0.539 1.202 0.756 0.623 0.544
ETTh2 0.353 0.395 0.341 0.373 0.517 0.491 0.461 0.451 0.528 0.505 0.456 0.460
ETTm1 0.679 0.520 0.465 0.425 0.740 0.557 0.520 0.464 0.880 0.615 0.584 0.500
LLM-based Models.
To further assess the generalization ability of the proposed PS loss, we extend our zero-shot forecasting experiments to LLM-based models, including OFA (Zhou et al., 2023), AutoTimes (Liu et al., 2024b), and Time-LLM (Jin et al., 2024), with the forecasting length set to 96. As shown in Table 12, incorporating PS loss consistently improves forecasting performance in most settings. These results demonstrate the compatibility of PS loss with LLM-based models and highlight its ability to enhance generalization in zero-shot forecasting.
Table 12: PS loss performance on LLM-based models for the zero-shot forecasting task. The backbones are OFA and Autotimes, with a forecasting length of 96.
Models OFA AutoTimes Time-LLM
Loss Function MSE loss PS loss MSE loss PS loss MSE loss PS loss
Source Target MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 ETTh2 0.289 0.347 0.292 0.343 0.303 0.362 0.305 0.363 0.287 0.349 0.286 0.346
ETTm1 0.723 0.532 0.742 0.534 0.728 0.552 0.724 0.549 0.747 0.566 0.706 0.538
ETTm2 0.219 0.311 0.216 0.303 0.231 0.323 0.233 0.325 0.225 0.312 0.224 0.310
ETTh2 ETTh1 0.413 0.419 0.412 0.419 0.469 0.470 0.418 0.432 0.487 0.465 0.422 0.425
ETTm1 0.762 0.540 0.742 0.534 0.987 0.618 0.842 0.573 0.814 0.574 0.706 0.533
ETTm2 0.209 0.302 0.208 0.299 0.248 0.327 0.236 0.322 0.216 0.308 0.222 0.308
ETTm1 ETTh1 0.495 0.474 0.493 0.472 0.541 0.497 0.527 0.488 0.530 0.484 0.502 0.468
ETTh2 0.331 0.376 0.333 0.377 0.347 0.395 0.324 0.378 0.320 0.372 0.316 0.368
ETTm2 0.179 0.262 0.180 0.262 0.190 0.276 0.185 0.271 0.195 0.276 0.197 0.277
ETTm2 ETTh1 0.511 0.485 0.469 0.460 0.544 0.508 0.500 0.477 0.549 0.497 0.471 0.453
ETTh2 0.306 0.364 0.297 0.351 0.298 0.362 0.294 0.354 0.311 0.369 0.294 0.352
ETTm1 0.411 0.408 0.359 0.380 0.445 0.429 0.373 0.390 0.411 0.416 0.357 0.378